What is the corresponding table for the fuctions x(x-2) and g(x)= 2x+4?
Answer by richard1234(7193)

First off, x(x-2) is not a well-defined function. For example, f(x) = x(x-2) and h(3x) = x(x-2) are functions.

Also, you cannot assign a unique "table" to a function (assuming you are referring to a table of values x, f(x)). The easiest way to construct a table is to pick some easy values of x and evaluate the function at that x.